# Bounds on the number of squares in recurrence sequences

### v1

@@ Abstract.
We investigate the number of squares in a very broad family of binary recurrence sequences with $u_{0}=1$. We show that there are at most two distinct squares in such sequences (the best possible result), except under such very special conditions where we prove there are at most three such squares.
@@ 1.1. Background
As with sequences of Lucas numbers, other binary recurrence sequences are important for the solution of Diophantine equations such as $aX^{2}-bY^{4}=c$, which is important in its own right, but is also a quartic model of elliptic curves.
@@ 1.2. Notation
To relate such sequences back to the quadratic equations mentioned in the previous section, observe that from (1.1),
@@ 1.3. Conjectures
If we replace $\varepsilon^{2k}$ in (1.1) with $\varepsilon^{k}$, it appears that the following is true.
@@
If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=p$ where $p$ is prime, or if $\left|N_{\alpha}\right|$ is a perfect square, then there are at most three distinct integer squares among the $y_{k}$’s.
We used PARI/GP [9] to search for squares among $y_{-60},\ldots,y_{60}$ from Conjecture 1.3, where $\varepsilon=\left(t+u\sqrt{d}\right)/2>1$ is the smallest algebraic integer that is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t,u\in\mathbb{Z}$ and $u\leq 10^{30}$. Note that this also includes searching among $y_{-30},\ldots,y_{30}$ from Conjectures 1.1 and 1.2. The range of $b$ and $d$ in this search was $1\leq b,d\leq 1000$ with $d$ squarefree. For $a$, we searched over two ranges: (1) $\max\left(1,\lfloor\sqrt{db^{4}}\rfloor-500\right)\leq a\leq\lceil\sqrt{db^{4}}\rceil+500$ (so $\left|N_{\alpha}\right|$ is small) and (2) $1\leq a\leq 1000$. For both ranges of $a$, we considered only $\gcd\left(a,b^{2}\right)$ squarefree. No counterexamples to any of these conjectures were found in this search.
The largest value of $|k|$ found in this search such that $y_{k}$ is a square was $k=-42$. One such example was found: $y_{-42}=12223^{2}$ for $(a,b,d,t,u)=(502,15,5,1,1)$. We also observed that no squares were found among the $y_{k}$’s with $|k|>30$ for $d>5$ and with $|k|>18$ for $d>13$.
@@ 1.5. Our method of proof
We have state and prove our results in Section 3 more generally than required here. This is because they will be useful in further work on binary recurrence sequences. An example is [18] where we prove Conjecture 1.2 when $N_{\alpha}$ is a prime power.
When $N_{\alpha}>0$ is not a square, we can apply Thue’s Fundamentaltheorem (see [14] and [15]) to obtain similar results.
@@ 1.6. Structure of this article
Section 2 concerns results on diophantine approximation, with a focus on the use of hypergeometric functions to do so. In Section 3, we collect various facts that we will require about elements of the sequence $\left(y_{k}\right)_{k=-\infty}^{\infty}$ defined by (1.1). These two sections are independent of each other. Their results are brought together in Section 4, where we state and prove Proposition 4.1. Our theorem follows from this proposition. Its proof is given in Section 5. Finally, in Section 6, we give examples showing that our results and conjectures are best possible.
@@ 2. Diophantine Approximation via Hypergeometric Functions
We can sometimes use the hypergeometric method to obtain effective irrationality measures that improve on Liouville’s result for the algebraic numbers that arise here. This was first done by Baker [2, 3] building on earlier applications of hypergeometric functions to diophantine approximation such as those of Siegel [11]. However, that does not suffice for us to prove our results here. The problem here arises not because of the exponent, $\mu$, in the effective irrationality measure, but because the constant, $c$, is too large. Upon investigating this further, we found that we can complete the proof of our results if we use not the effective irrationality measures from the hypergeometric method, but rather consider more carefully the actual results that we obtain from the use of hypergeometric functions.
@@ 2.1. Construction of Approximations
We let $D_{n,r}$ denote the smallest positive integer such that $D_{n,r}X_{m,n,r}(x)\in\mathbb{Z}[x]$ for all $m$ as above. For $d^{\prime}\in\mathbb{Z}$, we define $N_{d^{\prime},n,r}$ to be the largest integer such that $\left(D_{n,r}/N_{d^{\prime},n,r}\right)X_{m,n,r}\left(1-\sqrt{d^{\prime}}\,x\right)\in\mathbb{Z}\left[\sqrt{d^{\prime}}\right][x]$, again for all $m$ as above. We will use $v_{p}(x)$ to denote the largest power of a prime $p$ which divides into the rational number $x$. We put
@@
[…]
@@
However, the value $100$ in this inequality results in us requiring a lot of computation to complete the proof of our results here (in particular, Proposition 4.1). Therefore, we seek a smaller value at the expense of replacing $1.64$ by a larger value, whose value has less of an impact on our proof. For $r\geq 156$, $100\exp(1.64r)<0.2\exp(1.68r)$, so we compute directly the left-hand sides of (2.2) for $r\leq 155$. We find that the maximum values of the left-hand sides of (2.2) divided by $\exp(1.68r)$ both occur for $r=3$. Part (a) follows.
(b) We prove both of these by induction. A quick calculation shows that we have equality in both cases for $r=1$. Then we take the ratio of the expression for $r+1$ and divide by the expression for $r$. For the first inequality, this ratio is $(1/4)\left(r+5/4\right)/\left(r+3/2\right)$, so the inequality will hold if $\left(r+5/4\right)/\left(r+3/2\right)>\left(r/\left(r+1\right)\right)^{1/4}$. We take the fourth-power of both sides and subtract them, this gives us $\left(224r^{3}+944r^{2}+1329r+625\right)/\left(16\left(2r+3\right)^{4}\left(r+1\right)\right)$, which is clearly positive for $r\geq 1$. Since both $\left(r+5/4\right)/\left(r+3/2\right)$ and $\left(r/\left(r+1\right)\right)^{1/4}$ are positive real numbers for $r\geq 1$, the required inequality hlds.
@@
Put
$p_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega),$
$q_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}Y_{1,4,r}(\omega)\hskip 8.53581pt\text{ and}$
$R_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}R_{1,4,r}(\omega),$
where $r$ is any non-negative integer and $d^{\prime}$ will be determined below.
From Lemma 2.2(a), we have
$q_{r}\omega^{1/4}-p_{r}=R_{r}.$
By the definitions of $D_{4,r}$ and $N_{d^{\prime},4,r}$, we easily observe that $p_{r}^{\prime}$ and $q_{r}^{\prime}$ are algebraic integers of $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$. We can see that as follows. $X_{1,4,r}(z)$ is a polynomial of degree $r$ and
$p_{r}^{\prime}$ $=\frac{D_{1,4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega)$
$=\frac{D_{1,4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}\left(1-u_{2}\sqrt{t^{\prime}}\frac{-2}{u_{1}-u_{2}\sqrt{t^{\prime}}}\right).$
Since $d^{\prime}=u_{2}^{2}t^{\prime}$, $p_{r}^{\prime}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$. The analogous expression for $q_{r}^{\prime}$ shows that it is also an algebraic integer.
In fact, there may be further common factors. As in [16, Theorem 1], put
@@
Since $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/g$ is an algebraic integer, the argument above still applies to show that $p_{r}$ and $q_{r}$ are algebraic integers.
@@ 3.1. Representation Proposition
The proposition in this section plays a crucial role in this work. It is the expression for $f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ in this lemma that permits us to use the hypergeometric method described in the previous section. For this reason, we call it our representation proposition.
@@
where $\varepsilon=\left(t+u\sqrt{d}\right)/2\in\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t$ and $u$ non-zero rational integers, norm $N_{\varepsilon}$ and $\gcd\left(u,N_{\varepsilon}\right)=1$.
@@
[…]
@@
$f|\left(4b^{2}\operatorname{rad}\left(f^{\prime}\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right),$
where $f^{\prime}>0$, $f^{\prime}|\operatorname{sf}\left(N_{\alpha}\right)$ and $f^{\prime}<\sqrt{\operatorname{sf}\left(\left|N_{\alpha}\right|\right)}$, if $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)>1$ and $f^{\prime}=1$ otherwise.
(b) If $-N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha},N_{\varepsilon}\right)\right)\right)$. However, to do so, the first relationship in part (a) may become
$-f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
(c) If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=2^{\ell}p^{m}$ where $\ell,m\in\{0,1\}$ with $\ell+m\geq 1$ and $p$ is an odd prime, then we have $f\mid\left(4b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ when $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$, and $f\mid\left(2b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ otherwise.
@@
The condition that $\gcd\left(u,N_{\varepsilon}\right)=1$ is no restriction. If $g=\gcd\left(u,N_{\varepsilon}\right)$, then $g|t$. By equation (3.1), $g^{2}|x$ and $g^{2}|y^{2}$.
@@
Now suppose that $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$.
Put $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$. We shall prove the lemma in this case (i.e., $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$) for $r_{1}=r_{1}^{\prime}$. The proof is identical for $r_{1}=r_{1}^{\prime\prime}$.
If $k_{2}$ is even, then we have $2k_{1}<k_{2}$. If $k_{2}$ is odd, then we have $2k_{1}-1<k_{2}$. So in both cases, $2k_{1}\leq k_{2}$ (i.e., $v_{p}\left(4b^{2}\right)\leq v_{p}\left(r_{1}^{\prime 2}\right)$) and $v_{p}(y)\geq k_{1}$.
@@
Applying this, (3.4) and since $t^{2}+du^{2}=4N_{\varepsilon}+2du^{2}$, we obtain
@@
(3.7) $\min\left(v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right),v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right)\leq 0\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right).$
(c) If $r_{1}^{\prime}$ is even, then $v_{2}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<0\leq v_{2}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. The same result holds for $r_{1}^{\prime\prime}$.
@@
(a) We will show that for all primes, $p$, we have $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. From the definition of $\operatorname{rad}(\cdot)$, we have $v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)\leq 1$.
@@
(c) This follows immediately from the expression on the right-hand side of (3.8). ∎
@@
From the expressions for $r_{1}^{\prime}$ and $r_{1}^{\prime\prime}$, we need only show that $tb^{2}+au$ is even unless $N_{\alpha}\not\equiv 1\pmod{4}$ and $4\nmid d$. We do so by showing that if either (i) $tb^{2}$ even and $au$ odd or (ii) $tb^{2}$ odd and $au$ even, then $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
@@
Now suppose that $t$ is odd. Since $\left(t^{2}-du^{2}\right)/4\in\mathbb{Z}$ and $u$ is odd, it follows that $d\equiv 1\pmod{4}$. Here $b$ must be even, since $tb^{2}$ is even. Again $N_{\alpha}=a^{2}-db^{4}\equiv 1\pmod{4}$, since $a$ must be odd. Applying these to the right-hand side of (3.4), we find that $4y^{2}\equiv 2\pmod{4}$, so this case is not possible.
(ii) If $u$ is even, then since $t$ is odd, $N_{\varepsilon}$ cannot be an integer. So we must have $a$ even and $b$, $t$ and $u$ all odd. Since $t$ and $u$ are both odd, $d\equiv 1\pmod{4}$. So we have $N_{\alpha}\equiv 3\pmod{4}$. Hence $\left(b^{2}t+au\right)^{2}-N_{\alpha}u^{2}\equiv 2\pmod{4}$. But this quantity is $b^{2}$ times the right-hand side of (3.4), so it is $4b^{2}y^{2}$. Thus (ii) is not possible. ∎
@@
Applying (3.4), a routine calculation shows that with $r_{1}=tb^{2}+au\pm 2by$ and $s_{1}=-u$, we have
@@
We want to take the largest common factor, $g_{1}$, of $r_{1}$ and $s_{1}^{\prime}$ such that $4b^{2}r_{1}/g_{1}^{2}$ is also an integer. We have

### v2

@@ Abstract.
We investigate the number of squares in a very broad family of binary recurrence sequences with $u_{0}=1$. We show that there are at most two distinct squares in such sequences (the best possible result), except under very special conditions where we prove there are at most three such squares.
@@ 1.1. Background
As with sequences of Lucas numbers, other binary recurrence sequences are connected with the solutions of Diophantine equations such as $aX^{2}-bY^{4}=c$. Such equations are important in their own right, and also as quartic models of elliptic curves.
@@ 1.2. Notation
To relate such sequences back to the quartic equations mentioned in the previous section, observe that from (1.1),
@@ 1.3. Conjectures
In fact, a more general result than Conjecture 1.1 also appears to be true.
@@
Let $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be the sequence formed by replacing $\varepsilon^{2k}$ in (1.1) with $\varepsilon^{k}$.
@@
If $\left|N_{\alpha}\right|$ is a prime power or a perfect square, then there are at most three distinct integer squares among the $y_{k}$’s.
Examples showing that if these conjectures are true, then they are best possible are provided in Subsections 6.1, 6.2 and 6.3.
We used PARI/GP [10] to search for squares among $y_{-80},\ldots,y_{80}$ from Conjecture 1.3, where $\varepsilon=\left(t+u\sqrt{d}\right)/2>1$ is the fundamental unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t,u\in\mathbb{Z}$ and $u\leq 10^{30}$. Note that this also includes searching among $y_{-40},\ldots,y_{40}$ from Conjectures 1.1 and 1.2. The range of $b$ and $d$ in this search was $1\leq b\leq 2000$ and $1\leq d\leq 1000$ with $d$ squarefree. For $a$, we searched over two ranges: (1) $\max\left(1,\lfloor\sqrt{db^{4}}\rfloor-1000\right)\leq a\leq\lceil\sqrt{db^{4}}\rceil+1000$ (so $\left|N_{\alpha}\right|$ is small) and (2) $1\leq a\leq 2000$. For both ranges of $a$, we considered only $\gcd\left(a,b^{2}\right)$ squarefree. The search across these ranges took just under $319$ hours on a Windows 11 laptop with an Intel i7-13700H 2.40 GHz processor and 32GB of RAM. No counterexamples to any of these conjectures were found in this search.
@@ 1.4. Results
In Subsection 6.4, we present an infinite family of examples of sequences satisfying the conditions in part (a) with $y_{1}$ roughly one-tenth the size of the bounds in Proposition 4.1. The same is possible for $y_{-1}$ in part (a) and $y_{\pm 1}$ in part (b).
@@ 1.5. Our method of proof
We state and prove our results in Section 3 more generally than required here. This is because they will be useful in further work on binary recurrence sequences. An example is [18] where we prove Conjecture 1.2 when $N_{\alpha}$ is a prime power.
@@ 1.6. Structure of this article
Section 2 contains results on diophantine approximation, with a focus on the use of hypergeometric functions to do so. In Section 3, we collect various facts that we will require about elements of the sequence $\left(y_{k}\right)_{k=-\infty}^{\infty}$ defined by (1.1). These two sections are independent of each other. Their results are brought together in Section 4, where we state and prove Proposition 4.1. Our theorem follows from this proposition. Its proof is given in Section 5. Finally, in Section 6, we give examples showing that our results and conjectures are best possible.
1.7. Acknowledgements
The author deeply appreciates all the time and effort spent by Mihai Cipu, very carefully reading an earlier version of this work. His extensive comments, questions and our follow-up discussions led to significant improvements in the presentation of this work, as well as numerous corrections. The author is also grateful to the referee for their notes and suggestions, which improved this paper too.
@@ 2. Diophantine Approximation via Hypergeometric Functions
We can sometimes use the hypergeometric method to obtain effective irrationality measures that improve on Liouville’s result for the algebraic numbers that arise here. This was first done by Baker [2, 3] building on earlier applications of hypergeometric functions to diophantine approximation such as those of Siegel [12]. However, that does not suffice for us to prove our results here. The problem here arises not because of the exponent, $\mu$, in the effective irrationality measure, but because the constant, $c$, is too small. Upon investigating this further, we found that we can complete the proof of our results if we use not the effective irrationality measures from the hypergeometric method, but rather consider more carefully the actual results that we obtain from the use of hypergeometric functions.
@@ 2.1. Construction of Approximations
We let $D_{n,r}$ denote the smallest positive integer such that $D_{n,r}X_{m,n,r}(x)\in\mathbb{Z}[x]$ for all $m$ as above. For $d^{\prime}\in\mathbb{Z}$, we define $N_{d^{\prime},n,r}$ to be the largest integer such that $\left(D_{n,r}/N_{d^{\prime},n,r}\right)X_{m,n,r}\left(1-\sqrt{d^{\prime}}\,x\right)\in\mathbb{Z}\left[\sqrt{d^{\prime}}\right][x]$, again for all $m$ as above. We will use $v_{p}(x)$ to denote the largest power of a prime $p$ which divides the rational number $x$. We put
@@
(2.3) $\frac{5}{24\cdot 4^{r}r^{1/4}}\leq\frac{(1/4)\cdots(r+1/4)}{(r+1)\cdots(2r+1)}\quad\text{and}\quad\frac{r!\Gamma(3/4)}{\Gamma(r+3/4)}\leq 4r^{1/4}/3.$
@@
However, the value $100$ in this inequality results in us requiring a lot of computation to complete the proof of our results here (in particular, Proposition 4.1). Therefore, we seek a smaller value at the expense of replacing $1.64$ by a larger value, whose value has less of an impact on our proof. For $r\geq 156$, we have $100\exp(1.64r)<0.2\exp(1.68r)$, so we compute directly the left-hand sides of (2.2) for $r\leq 155$. We find that the maximum values of the left-hand sides of (2.2) divided by $\exp(1.68r)$ both occur for $r=3$. Part (a) follows.
(b) We prove both of these by induction. A quick calculation shows that we have equality in both cases for $r=1$. Then we take the expression for $r+1$ and divide it by the expression for $r$. For the first inequality, this ratio is $(1/4)\left(r+5/4\right)/\left(r+3/2\right)$, so the inequality will hold if $\left(r+5/4\right)/\left(r+3/2\right)>\left(r/\left(r+1\right)\right)^{1/4}$. We take the fourth-power of both sides and subtract them, this gives us
$\frac{224r^{3}+944r^{2}+1329r+625}{16\left(2r+3\right)^{4}\left(r+1\right)},$
which is clearly positive for $r\geq 1$. Since both $\left(r+5/4\right)/\left(r+3/2\right)$ and $\left(r/\left(r+1\right)\right)^{1/4}$ are positive real numbers for $r\geq 1$, the required inequality hlds.
@@
As in [16, Theorem 1], put
@@
By the definitions of $D_{4,r}$ and $N_{d^{\prime},4,r}$, we see that $X_{1,4,r}(1-z)$ is a polynomial of degree $r$ with coefficients in $\mathbb{Z}\left[\sqrt{d^{\prime}}\right]$ and
$p_{r}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega)$
$=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}\left(1-u_{2}\sqrt{t^{\prime}}\frac{-2}{u_{1}-u_{2}\sqrt{t^{\prime}}}\right).$
Since $d^{\prime}=u_{2}^{2}t^{\prime}$ and $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/(2g)$ is an algebraic integer, it follows that $p_{r}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$. The analogous expression for $q_{r}$ shows that it is also an algebraic integer.
@@ 3.1. Representation Proposition
The proposition in this section plays a crucial role in this work. It is the expression for $f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ in this proposition that permits us to use the hypergeometric method described in the previous section. For this reason, we call it our representation proposition.
@@
where $\varepsilon=\left(t+u\sqrt{d}\right)/2\in\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t$ and $u$ non-zero rational integers, norm $N_{\varepsilon}$.
@@
$f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\quad\text{and}$
@@
$f|\left(4b^{2}\operatorname{rad}\left(f^{\prime}\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right),$
where $f^{\prime}|\operatorname{sf}\left(N_{\alpha}\right)$ and $0<f^{\prime}<\max\left(2,\sqrt{\left|\operatorname{sf}\left(N_{\alpha}\right)\right|}\right)$.
(b) If $-N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, where the first relationship in part (a) is replaced by
±$\pm f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
(c) If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=2^{\ell}p^{m}$ where $\ell,m\in\{0,1\}$ with $\ell+m\geq 1$ and $p$ is an odd prime, then we have $f\mid\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ when $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$, and $f\mid\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ otherwise.
@@
(2) Now suppose that $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$.
Put $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$. We shall prove the lemma in this case (i.e., when $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$) for $r_{1}=r_{1}^{\prime}$. The proof is identical for $r_{1}=r_{1}^{\prime\prime}$.
If $k_{2}$ is even, then we have $2k_{1}<k_{2}$. If $k_{2}$ is odd, then we have $2k_{1}-1<k_{2}$. So in both cases, $2k_{1}\leq k_{2}$ (i.e., $v_{p}\left(4b^{2}\right)\leq v_{p}\left(r_{1}^{\prime}\right)$) and $v_{p}(y)\geq k_{1}$.
@@
Applying this and the right-hand expression in (3.4) for $4y^{2}$, we obtain
@@
(3.7) $\min\left(v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right),v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right)\leq 0.$
(c) If $r_{1}^{\prime}$ is even, then $v_{2}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<0$. The same result holds for $r_{1}^{\prime\prime}$.
@@
(a) We will prove that for all primes, $p$, we have $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. The proof is the same for $r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}$.
@@
(c) This follows from (3.8). We have $v_{2}\left(\operatorname{sf}\left(r_{1}^{\prime}\right)\right)=0$ or $1$. In both cases, since $r_{1}^{\prime}$ is even, we find that $v_{2}$ of the denominator of the right-hand side of (3.8) is strictly greater than $v_{2}$ of its numerator. ∎
@@
From the expressions for $r_{1}^{\prime}$ and $r_{1}^{\prime\prime}$, we need only show that $tb^{2}+au$ is even unless $N_{\alpha}\equiv 1\pmod{4}$ and $4\mid d$. We do so by showing that if either (i) $tb^{2}$ even and $au$ odd or (ii) $tb^{2}$ odd and $au$ even, then $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
@@
Now suppose that $t$ is odd. Since $\left(t^{2}-du^{2}\right)/4\in\mathbb{Z}$ and $u$ is odd, it follows that $d\equiv 1\pmod{4}$. Here $b$ must be even, since $tb^{2}$ is even. Again $N_{\alpha}=a^{2}-db^{4}\equiv 1\pmod{4}$, since $a$ must be odd. Applying these to the middle expression in (3.4), we find that $4y^{2}\equiv 2\pmod{4}$, so this case is not possible.
(ii) If $u$ is even, then since $t$ is odd, $N_{\varepsilon}$ cannot be an integer. So we must have $a$ even and $b$, $t$ and $u$ all odd. Since $t$ and $u$ are both odd, $d\equiv 1\pmod{4}$. So we have $N_{\alpha}\equiv 3\pmod{4}$. Hence $\left(b^{2}t+au\right)^{2}-N_{\alpha}u^{2}\equiv 2\pmod{4}$. But this quantity is $b^{2}$ times the middle expression in (3.4), so it is $4b^{2}y^{2}$. Thus (ii) is not possible. ∎
@@
With $r_{1}=tb^{2}+au\pm 2by$ and $s_{1}=-u$, a routine calculation, along with (3.4), shows that
@@
We want to take the largest common factor, $g_{1}$, of $r_{1}$ and $s_{1}^{\prime}$ such that $4b^{2}r_{1}/g_{1}^{2}$ is also an integer. That is,